OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (theeric):

Hi? Are you there?

OpenStudy (theeric):

Here will be references for you: You need to know slope-intercept form. It is that one form where you can easily see slope and \(y\)-intercept. It is \(y=mx+b\) where \(m\) is the slope and \(b\) is the \(y\)-intercept. A line that is perpendicular to another line has a negative reciprocal slope of the other. How do you know the slope of the other? It is in slope-intercept form, so you can look at what must be \(m\); what is multiplying \(x\). So, you can find your slope that way. Then you can use point-slope formula, \(\dfrac{y-y_1}{x-x_1}=m\). You know that your point on the line is \((x_1,\ y_1)\) and you know the slope, \(m\), from finding the negative reciprocal of the other slope. Then solve for \(y\) in that last equation to see the slope-intercept form.